OpenStudy (anonymous):

NEED SOMEONE TO RECHECK ANS *****HELP PLEASE If P (A) = 0.2 and P (B) = 0.3 and A and B are disjoint, what is P (A or B)? (1 point) 0.00 0.06 0.10 0.44 *0.50 4. Given that you roll a fair six-sided die, what is the probability that you roll an odd number? (1 point) 0 *1/6 1/3 1/2 1 5. There are 3 red marbles, 4 white marbles, and 1 green marble in a bag. Marbles are drawn without replacement. What is the probability that 3 marbles can be drawn without drawing the green marble? (1 point) *0.625 0.643 0.670 0.998 0.179 6. Which of the following represents the sample space for rolling a six-sided die and tossing a coin? (1 point) {1H, 2H, 3H, 4H, 5H, 6H, 1T, 2T, 3T, 4T, 5T, 6T} *{1, 2, 3, 4, 5, 6} {H, T} {1, 2, 3, 4, 5, 6, H, T} 7. If P (A) = 0.2 and P (B) = 0.3 and A and B are independent but NOT necessarily disjoint, find P (A or B). (1 point) 0.06 0.10 *0.44 0.50 Cannot be determined from the information given 8. If three coins are tossed, what is the number of equally likely outcomes? (1 point) 3 4 *6 8 9

OpenStudy (anonymous):

3. You'll be using this formula: \[P(A\cup B)=P(A)+P(B)-P(A\cap B)\] 4. How many odd numbers can you possible roll? How many possible events are there? The probability would be \[\frac{\text{possible odds}}{\text{total possible}}\] 5. You're looking for \(P(\text{not green})\). It would help to draw yourself a probability tree. 6. This one should be fairly obvious. You can immediately eliminate the second and third options because each considers only the die, or only the coin. The experiment involves rolling a die AND tossing a coin, and, presumably, you would be observing both events happening at the same time, so you can also eliminate the fourth option. 7. Same formula as for (3), with the fact that if \(A\) and \(B\) are independent, you have \(P(A\cap B)=P(A)P(B)\). 8. List the sample space, then find the probability of each event. Count the number of outcomes with equal probability.

OpenStudy (anonymous):

3.AUB=0.2+0.3-(0.2)(0.3)=0.44

OpenStudy (anonymous):

4. POSSIBLE ODDS/TOTAL POSSIBLE=1/6

OpenStudy (anonymous):

Your answer for 3 is not correct. That is the answer for 7. 4 is correct.

OpenStudy (anonymous):

so how do I do it

OpenStudy (anonymous):

I use the formula that I was given

OpenStudy (anonymous):

so number 3 is 50 6. 1,2,3,4,5, and 6

OpenStudy (anonymous):

7. 44 8.6